OpenStudy (anonymous):

Write an expression to represent:Three less than the product of two and a number x

OpenStudy (amilapsn):

Product of 2 and a number x: \(2\times x\) Three less than the product of two and that x: \(3-2\times x\)

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Ummm since it is supposed to be three less... I think it should be (x2)-3

OpenStudy (anonymous):

idk

OpenStudy (♪chibiterasu):

Ok so lets break it down. Two = 2 Numer x = x Three Less than = 3 the product of two and number x: (2x) Less than three = - 3 Combine: (2x) - 3

OpenStudy (♪chibiterasu):

Product = multiplication Less than = subraction

OpenStudy (anonymous):

soo (2x)-3?

OpenStudy (amilapsn):

Oh.... I got it wrong sorry @jamesr :(

OpenStudy (anonymous):

u gud we all make mistakes. @amilapsn

OpenStudy (anonymous):

@♪Chibiterasu is that it

OpenStudy (amilapsn):

@♪Chibiterasu if there's not "than" then my answer is correct right?

OpenStudy (♪chibiterasu):

No. "the product of two and a number x" is a separate value. Other wise it would be "three less than two times a number x".

OpenStudy (♪chibiterasu):

So (2x) - 3 would be right.

OpenStudy (amilapsn):

3 less 4=3-4, right @♪Chibiterasu ?